OpenStudy (anonymous):

Five times the sum of a number and 27 is greater than or equal to six times the sum of that number and 26. What is the solution set of this problem?

OpenStudy (anonymous):

Options A. (−∞,−21) B.(−∞,−21] C.[-21,+∞) D.(21,+∞)

OpenStudy (anonymous):

gotcha, i read the question wrong 5(x+27) >= 6(x+26) 5x + 135 >= 6x +156 -21 >= x or x <= -21

OpenStudy (anonymous):

thank you!! (: